Hydrogen peroxide is a powerful oxidizing agent that is used in concentrated solution in rocket fuels and dilute solution as a hair bleach.

A 30% by mass of H2O2 solution has a density of 1.11g/ml. Calculate it's morality, mole fraction of H2O2, and molarity.


1
Expert's answer
2021-04-02T05:12:05-0400

soln = solution

The molar mass of H2O2 is 34 g mol-1.

The molar mass of H2O is 18 g mol-1.


Solution:

Assume that 100 g solution is given.


1) Determine the molarity of the solution:

(1.11 g soln / mL soln) × (30 g H2O2 / 100 g soln) × (1 mol H2O2 / 34 g H2O2) × (1000 mL / 1L) = 9.79 M

The molarity of H2O2 solution is 9.79 M


2) Determine the mole fraction of H2O2:

χ(H2O2) = Moles of H2O2 / Total moles of solution


Moles of H2O2 = (30 g H2O2) × (1 mol H2O2 / 34 g H2O2) = 0.88235 mol

Moles of H2O = (70 g H2O) × (1 mol H2O / 18 g H2O) = 3.8889 mol

Total moles of solution = Moles of H2O2 + Moles of H2O = 0.88235 mol + 3.8889 mol = 4.77125 mol


χ(H2O2) = 0.88235 mol / 4.77125 mol = 0.18493 = 0.185

The mole fraction of H2O2 is 0.185


Answer:

The molarity of H2O2 solution is 9.79 M

The mole fraction of H2O2 is 0.185